# Supernova Remnants with M dwarf surviving companions

### v1

@@ 1 Introduction
The evolutionary scenario first proposed to produce a Type Ia supernova from a WD (Whelan & Iben, 1973; Nomoto, 1982), called the single-degenerate scenario (SD), assumes the growth in mass of a WD up to the Chandrasekhar mass by accretion of material from a non-degenerate companion through Roche-lobe overflow. Rapid contraction then explosively ignites the central layers of the WD (made of C+O) when reaching a critical density. The explosive thermonuclear burning would start as a deflagration, or a deflagration that evolves into a detonation farther from the center (delayed detonation). The companion, although hit by the SN ejecta, can survive (Marietta et al., 2000; Pan et al., 2010, 2012a; Liu et al., 2012; Rau & Pan, 2022).
@@
In the SD scenario, the companion star of the exploding WD is ejected with the orbital velocity plus the kick due to the impact of the SN material. Velocities in the hundreds of km s-1 should be expected in this case (Pan et al., 2012a). Post-explosion evolution of different possible SN Ia companions has been calculated by Pan et al. (2012a, 2014); Rau & Pan (2022).
@@
On the contrary, if the companion star were another WD (Webbink, 1984; Iben & Tutukov, 1984), which is called the double-degenerate (DD) scenario, the companion should be destroyed by the mass-transfer process itself. In this model, the orbit of the two WDs decreases due to the emission of gravitational waves. Eventually, tidal interaction disrupts the companion (the less massive WD or secondary) and its debris merge with the primary WD, increasing its mass towards the Chandrasekhar limit. Recent modelling, however, has shown that the explosion can be produced differently. So, in the “violent merging” model, the interaction of the debris of the secondary WD after its collision with the primary can trigger the explosion (Pakmor et al., 2010, 2012; Sato et al., 2015).
@@
In the core-degenerate (CD) model, a WD merges with the electron-degenerate core of a red-giant or AGB star (Soker, 2013; Soker et al., 2014; Soker, 2019). In this case, the WD is engulfed by the envelope of the companion, and friction leads to merging. If there is sufficient delay between merging and explosion (the merger to explosion delay, MED, model; Soker 2019, 2022), the exploding object may have reached the Chandrasekhar mass and undergo central ignition, developing into a delayed detonation. No companion is left in this scenario.
@@
There is a new turn on the outcome of the DD scenario: the so-called Dynamically-Driven Double-Degenerate Double-Detonation model (D6; see Shen 2025, for instance, for a recent account), where the explosion happens when the less massive WD has shed only a part of its mass, the rest being still gravitationally bound. If the secondary were a He WD, a He detonation could be produced by the He accumulated, by steady mass transfer, on top of the primary (made of C+O). The ensuing compression can induce a second detonation close to the center. Due to the extreme closeness of the two WDs at the time of the SN Ia explosion, in this model, the surviving WD should be ejected at very high velocities. This can be the origin of hypervelocity stars (WDs with velocities km s-1) with peculiar characteristics. It has also been proposed that high-velocity stars could result from the only partial burning of a WD in a SN Iax explosion (Raddi et al., 2019).
@@
A possible surviving companion of a SN Ia has recently been discovered in the Galactic supernova remnant (SNR) G272.2-3.2 (Ruiz-Lapuente et al., 2023), using the data from the Gaia satellite. It has been named MV-G272. The star is a 8.9 outlier in proper motion, and its trajectory places it at the center of the remnant at the time of the SN explosion, 7500 yr ago. MV-G272 is a M1-M2 dwarf with a large space velocity and is slowly rotating. Here, we investigate the way a M dwarf star may show the characteristics of MV-G272 (space and rotational velocities, mass, luminosity, effective temperature, and surface gravity) after being hit by the ejecta of a SN Ia and has been cooling down for the SNR age afterwards. In this paper, we investigate in depth the outcome of SNe Ia explosions with low-mass main-sequence companions. We explore the evolution with time after the explosion in the H-R diagram. We determine if MV-G272 can be the companion of the SNIa that gave rise to the remnant G272.2-3.2.
@@ 2 Observations
As mentioned in the Introduction, the hypervelocity or high-velocity stars proposed as surviving companions of SNe Ia do not show fast rotations either. Three stars have their rotational velocities measured (LP 398-9, LP 40-365, and J0546+0836; see Table 2). In LP 398-9, a periodic 15.4 h signal in the UV and optical light curves has been detected by Chandra et al. (2022) and interpreted as surface rotation. LP 40-365 has been attributed to a partially burned WD in a Type Iax SN. A photometric variation corresponding to a 5.8% in the UV flux, with a period of 8.914 m has been measured by Hermes et al. (2021). The rotational velocity of J0546+0836 has been estimated by El-Badry et al. (2023), from the broadening of the 5801, 5012 doublet of C IV. In Table 2, their periods, radii, linear and rotational velocities, together with those of MV-G272, as well as the proportion of the rotational to linear velocities, expressed in percentages, are shown.
| Object | sin | % | |||
| [hr] | [] | [km s-1] | [km s-1] | ||
| MV-G272 | 66.92 | 0.564 | 363 | 4 | 1.1 |
| D6-1 | 2045 | 20 | 0.97 | ||
| LP 398-9 (D6-2) | 15.4 | 0.200.01 | 1151 | 15.78 | 1.37 |
| LP 40-365 (GD 492) | 8.914 | 0.160.01 | 837 | 21.80 | 2.6 |
| J0546+0836 | 0.33∗ | 0.051 | 1699 | 180 | 10.6 |
@@ 3.1 Companion Models
We use the stellar evolution code MESA, version r10398 (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023), to construct three low-mass companion models with masses , and and with metallicity . Since the detailed binary evolution in the single-degenerate scenario prior to SN Ia explosion remains elusive, we assume that these companion models have the same structure as at their Zero-Age Main Sequence (ZAMS) stage because their MS lifetimes are much longer than the delayed time of a SN Ia, and we neglect binary interactions during the binary evolution phase for simplicity. The corresponding stellar radii of these companion models are , , and . Figure 1 shows the density profiles of these three companion models.
@@ 3.2 Supernova Impact
In three-dimensional simulations, we additionally include symmetric-breaking effects, such as the spin and orbital motions of the companion star. The orbital velocity is calculated using Kepler’s law, assuming zero eccentricity. A spin-to-orbital ratio of is assumed in this study. We also ensure that the region between the center of the SN explosion and the companion star remains at the highest refinement level. The SN ejecta is approximated as pure nickel (56Ni) and used as a tracer in simulations.
Table 3 summarizes all the 2D and 3D models we considered in the hydrodynamics simulations. Note that in the standard single-degenerate scenario, the binary companions are expected to experience Roche-lobe overflow at the onset of the SN explosion. The binary separations are typically around three times the companion’s radius. In this study, the binary separations are fixed at approximately , corresponding to , which falls within the range of the standard SDS.
@@ 3.4 Stellar rotation and magnetic braking
We incorporate magnetic braking into our models, assuming a magnetic field strength of . Note that the high magnetic fields, such as or 5000 G, have been invoked to explain the orbital decays of black hole X-ray binaries (González Hernández et al., 2014, 2017). The magnetic braking implementation follows the simple angular momentum loss formula described in Weber & Davis (1967), as provided in the standard test_suites of MESA. Magnetic braking is activated whenever a mass loss occurs during the simulation. Finally, we investigate the differences in the evolution with respect to different values of and .
@@ 4 Results
In this section, we present the results of our multidimensional hydrodynamic simulations on the impact of supernova ejecta on a low-mass companion. We perform ten two-dimensional simulations using three low-mass companion models with masses of 0.5, 0.6, or 0.7 . For each companion model, we explore different initial binary separations . Table 3 summarizes the parameter space explored in this study and lists a few important physical quantities from the simulation outcomes. We select four representative cases from the two-dimensional simulations and conduct the corresponding three-dimensional simulations, including spin and orbital motions. In addition, we perform three additional three-dimensional simulations without spin and orbital motions as reference simulations (see Table 3).
@@ 4.2 Angular momentum loss
Figure 5 shows the evolution of the companion’s spin angular momentum (magnitude) as a function of time for several representative 3D models that include the initial rotation. We calculate the spin angular momentum of the bound remnant (excluding unbound ejecta) in the companion’s center-of-mass frame. It is evident that the SN impact causes a net loss of angular momentum from the star. As the supernova ejecta sweeps over the companion, it strips away a portion of the star’s outer layers and, with them, carries off a substantial fraction of the star’s angular momentum. This is a key channel for the spin-down of a surviving companion as reported in Pan et al. (2012a).
@@ 4.3 Post-impact evolution
Figure 9 shows the evolution of the companion’s surface rotation velocity under the influence of magnetic braking for various assumed magnetic field strengths. As expected, higher initial field strengths lead to more efficient angular momentum loss via the stellar wind and, thus, a faster spin-down of the star. In the absence of magnetic braking, the post-impact contraction would cause the surface rotation velocity () to rise (due to the shrinking moment of inertia). However, with sufficiently strong magnetic torques, this spin-up can be counteracted and reversed. Our simulations indicate that extraordinarily strong surface fields on the order of kG (kilogauss) are required to reduce the rotation to km s-1 within a practical observational timeframe after the explosion. For example, with kG, the companion’s rotation is efficiently braked to a slow spin by the time the SN remnant is a few thousand years old. We have assumed a constant stellar wind mass-loss rate of yr-1 in these post-impact evolution calculations, consistent with a robust magnetized wind from a late-type star. We note that for weaker field strengths (significantly below a few kG), the magnetic braking is too feeble to offset the contraction-induced spin-up, and the survivor remains a relatively fast rotator. Thus, while the hydrodynamic impact itself removes a large fraction of the companion’s original angular momentum (Section 4.2), an additional mechanism—such as a very strong magnetic wind—is likely necessary to brake the star’s rotation to the low values that might be observed in candidate SN companion stars.

### v2

# Supernova Ia Remnants with M dwarf surviving companions

@@ 1 Introduction
The evolutionary scenario first proposed to produce a Type Ia supernova from a WD (Whelan & Iben, 1973; Nomoto, 1982), called the single-degenerate scenario (SD), assumes the growth in mass of a WD up to the Chandrasekhar mass by accretion of material from a non-degenerate companion through Roche-lobe overflow. Rapid contraction then explosively ignites the central layers of the WD (made of C+O) when reaching a critical density. The explosive thermonuclear burning would start as a deflagration, or a deflagration that evolves into a detonation farther from the center (delayed detonation). The companion, although hit by the SN ejecta, can survive (Marietta et al., 2000; Pan et al., 2010, 2012a; Liu et al., 2012; Rau & Pan, 2022; Liu et al., 2023).
@@
In the SD scenario, the companion star of the exploding WD is ejected with the orbital velocity plus the kick due to the impact of the SN material. Velocities in the hundreds of km s-1 should be expected in this case (Pan et al., 2012a). Post-explosion evolution of different possible SN Ia companions has been calculated by Wheeler et al. (1975); Pan et al. (2012a, 2014); Bauer et al. (2019); Liu et al. (2021); Rau & Pan (2022); Wong et al. (2024); Wang et al. (2025).
@@
On the contrary, if the companion star were another WD (Webbink, 1984; Iben & Tutukov, 1984), which is called the double-degenerate (DD) scenario, the companion should be destroyed by the mass-transfer process itself. In this model, the orbit of the two WDs decreases due to the emission of gravitational waves. Eventually, tidal interaction disrupts the companion (the less massive WD or secondary) and its debris merges with the primary WD, increasing its mass towards the Chandrasekhar limit. Recent modelling, however, has shown that the explosion can be produced differently in the “violent merger” DD scenario, the interaction of the debris of the secondary WD after its collision with the primary can trigger a prompt detonation while the merger is still ongoing, given as a result a SN explosion whose characteristics have been seen in some observed SN Ia (Pakmor et al., 2010, 2012; Sato et al., 2015).
@@
In the core-degenerate (CD) model, a WD merges with the electron-degenerate core of an AGB star (Soker, 2013; Soker et al., 2014; Soker, 2019). In this case, the WD is engulfed by the envelope of the companion, and friction leads to merging. If there is sufficient delay between merging and explosion (the merger to explosion delay, MED, model; Soker 2019, 2022), the exploding object may have reached the Chandrasekhar mass and undergo central ignition, developing into a delayed detonation.
@@
There is a new turn on the outcome of the DD scenario: the so-called Dynamically-Driven Double-Degenerate Double-Detonation model (D6; see Shen 2025, for instance, for a recent account), where the explosion happens when the less massive WD has shed only a part of its mass, the rest being still gravitationally bound. If the secondary were a He WD, a He shell detonation could be triggered by the violent impact of the accretion stream on the primary WD. The ensuing compression can induce a second detonation close to the center. Due to the extreme closeness of the two WDs at the time of the SN Ia explosion, in this model, the surviving WD should be ejected at very high velocities. This can be the origin of hypervelocity stars (WDs with velocities km s-1) with peculiar characteristics. It has also been proposed that high-velocity stars could result from the only partial burning of a WD in a SN Iax explosion (Raddi et al., 2019).
@@
A possible surviving companion of a SN Ia has recently been discovered in the Galactic supernova remnant (SNR) G272.2-3.2 (Ruiz-Lapuente et al., 2023), using the data from the Gaia satellite. It has been named MV-G272. The star is a 8.9 outlier in proper motion, and its trajectory places it at the center of the remnant at the time of the SN explosion, 7500 yr ago. MV-G272 is a M1-M2 dwarf with a large space velocity and is slowly rotating. Here, we investigate the way a M dwarf star may show the characteristics of MV-G272 (space and rotational velocities, mass, luminosity, effective temperature, and surface gravity) after being hit by the ejecta of a SN Ia and has been cooling down for the SNR age afterwards.
In the standard single-degenerate (SD) scenario for SNe Ia, M dwarfs are generally not considered likely binary companions due to their long delay time before initiating stable Roche-lobe overflow. However, Wheeler (2012) proposed a novel SD channel in which WD+M-dwarf binaries could produce SNe Ia if they emerge from a common-envelope phase and mass transfer is aided by the “magnetic bottle” effect of a fully convective, low-mass M-dwarf companion. In this paper, we investigate in depth the outcome of SNe Ia explosions with low-mass main-sequence companions. We explore the evolution with time after the explosion in the H-R diagram. We determine if MV-G272 can be the companion of the SNIa that gave rise to the remnant G272.2-3.2.
@@ 2 Observations
As mentioned in the Introduction, the hypervelocity or high-velocity stars proposed as surviving companion candidates of SNe Ia do not show fast rotations either. Three stars have their rotational velocities measured (LP 398-9, LP 40-365, and J0546+0836; see Table 2). In LP 398-9, a periodic 15.4 h signal in the UV and optical light curves has been detected by Chandra et al. (2022) and interpreted as surface rotation. LP 40-365 has been attributed to a partially burned WD in a Type Iax SN. A photometric variation corresponding to a 5.8% in the UV flux, with a period of 8.914 h has been measured by Hermes et al. (2021). The rotational velocity of J0546+0836 has been estimated by El-Badry et al. (2023), from the broadening of the 5801, 5012 doublet of C IV. In Table 2, their periods, radii, linear and rotational velocities, together with those of MV-G272, as well as the proportion of the rotational to linear velocities, expressed in percentages, are shown.
| Object | sin | % | Ref. | |||
| [hr] | [] | [km s-1] | [km s-1] | |||
| MV-G272 | 66.92 | 0.564 | 363 | 4 | 1.1 | Ruiz-Lapuente et al. (2023) |
| D6-1 | 2045 | 20 | 0.97 | Shen (2025), private communication | ||
| LP 398-9 (D6-2) | 15.4 | 0.200.01 | 1151 | 15.78 | 1.37 | Chandra et al. (2022) |
| LP 40-365 (GD 492) | 8.914 | 0.160.01 | 837 | 21.80 | 2.6 | Hermes et al. (2021) |
| J0546+0836 | 0.33∗ | 0.051 | 1699 | 180 | 10.6 | El-Badry et al. (2023) |
@@ 3.1 Companion Models
We use the stellar evolution code MESA, version r10398 (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023), to construct three low-mass companion models with masses , and and with metallicity . Since the detailed binary evolution in the single-degenerate scenario prior to a SN Ia explosion remains elusive, it is possible that the binary system underwent a common-envelope stage, leading to a low-mass binary companion prior to the explosion (Wheeler, 2012). We assume that these companion models have the same structure as at their Zero-Age Main Sequence (ZAMS) stage because their MS lifetimes are much longer than the delayed time of a SN Ia, and we neglect binary interactions during the binary evolution phase for simplicity. The corresponding stellar radii of these companion models are , , and . Figure 1 shows the density profiles of these three companion models.
@@ 3.2 Supernova Impact
In three-dimensional simulations, we additionally include symmetric-breaking effects, such as the spin and orbital motions of the companion star. The orbital velocity is calculated using Kepler’s law, assuming zero eccentricity. A spin-to-orbital ratio of is assumed in this study. We also ensure that the region between the center of the SN explosion and the companion star remains at the highest refinement level. The composition of the SN ejecta is approximated as pure nickel (56Ni) and is used as a tracer in simulations (Pan et al., 2012a).
Table 3 summarizes all the 2D and 3D models we considered in the hydrodynamics simulations. Note that in the standard single-degenerate scenario, the binary companions are expected to experience Roche-lobe overflow at the onset of the SN explosion. The binary separations are typically around three times the companion’s radius. In addition to the standard SD scenario, we consider extra systems that may have exited Roche-lobe overflow prior to explosion, for instance, due to prior mass transfer, a common-envelope phase, or spin-up/spin-down delays. For this reason, the binary separations in our models are fixed at approximately , corresponding to , which intentionally extend beyond the canonical Roche-lobe overflow range.
@@ 3.4 Stellar rotation and magnetic braking
The surface magnetic fields of typical inactive and slowly rotating M dwarf stars are usually about a few hundred gauss (Reiners & Basri, 2007), but can reach several thousands gauss in fast-rotating stars (Reiners et al., 2009; Shulyak et al., 2017). It is also unclear whether the magnetic fields would be destroyed or amplified during the SN impact. In this study, we incorporate magnetic braking into our models, assuming a surface magnetic field strength of . Note that the high magnetic fields, such as or 5000 G, have been invoked to explain the orbital decays of black hole X-ray binaries (González Hernández et al., 2014, 2017). The magnetic braking implementation follows the simple angular momentum loss formula described in Weber & Davis (1967), as provided in the standard test_suites of MESA. Magnetic braking is activated whenever a mass loss occurs during the simulation. Finally, we investigate the differences in the evolution with respect to different values of and .
@@ 4 Results
, we present the results of our multidimensional hydrodynamic simulations on the impact of supernova ejecta on a low-mass companion. We perform ten two-dimensional simulations using three low-mass companion models with masses of 0.5, 0.6, or 0.7 . For each companion model, we explore different initial binary separations . Table 3 summarizes the parameter space explored in this study and lists a few important physical quantities from the simulation outcomes. We select four representative cases from the two-dimensional simulations and conduct the corresponding three-dimensional simulations, including spin and orbital motions. In addition, we perform three additional three-dimensional simulations without spin and orbital motions as reference simulations (see Table 3).
@@ 4.2 Angular momentum loss
Figure 5 shows the evolution of the companion’s spin angular momentum (magnitude) as a function of time for several representative 3D models that include the initial rotation. We calculate the spin angular momentum of the bound remnant (excluding unbound ejecta) in the companion’s center-of-mass frame. It is evident that the SN impact causes a net loss of angular momentum from the star. As the supernova ejecta sweeps over the companion, it strips away a portion of the star’s outer layers and, with them, carries off a substantial fraction of the star’s angular momentum. This is a key channel for the spin-down of a surviving companion as reported in Pan et al. (2012a); Liu et al. (2013).
@@ 4.3 Post-impact evolution
Figure 9 shows the evolution of the companion’s surface rotation velocity under the influence of magnetic braking for various assumed magnetic field strengths. As expected, higher initial field strengths lead to more efficient angular momentum loss via the stellar wind and, thus, a faster spin-down of the star. In the absence of magnetic braking, the post-impact contraction would cause the surface rotation velocity () to rise (due to the shrinking moment of inertia). However, with sufficiently strong magnetic torques, this spin-up can be counteracted and reversed. Our simulations indicate that extraordinarily strong surface fields on the order of kG (kilogauss) are required to reduce the rotation to km s-1 within a practical observational timeframe after the explosion. For example, with kG, the companion’s rotation is efficiently braked to a slow spin by the time the SN remnant is a few thousand years old. We have assumed a constant stellar wind mass-loss rate of yr-1 in these post-impact evolution calculations, consistent with a robust magnetized wind from a late-type star. We note that for weaker field strengths (significantly below a few kG), the magnetic braking is too feeble to offset the contraction-induced spin-up, and the survivor remains a relatively fast rotator. Thus, while the hydrodynamic impact itself removes a large fraction of the companion’s original angular momentum (Section 4.2), an additional mechanism—such as a very strong magnetic wind—is likely necessary to brake the star’s rotation to the low values that might be observed in candidate SN companion stars as suggested by Ruiz-Lapuente et al. (2023).